at noon, a car travels from A to B at a uniform speed of 60 kilometers per hour. At the same time, a van travels uniformly from B to A. At 3:30 p.m., the car and van are 80 kilometers apart after having passed each other earlier.If the car arrives at B at 5 p.m., find the time at which the van will arrive at A.
Answer by ankor@dixie-net.com(22740)

t noon, a car travels from A to B at a uniform speed of 60 kilometers per hour.
At the same time, a van travels uniformly from B to A.
At 3:30 p.m., the car and van are 80 kilometers apart after having passed each other earlier.
If the car arrives at B at 5 p.m., find the time at which the van will arrive at A.
;
Find the A to B distance from the information given
5 * 60 = 300 km
:
Find how far the car is from A when they are 80 km apart
3.5 * 60 = 210 km
:
Find how far from A is the Van at 3:30 when the car is 210 km from A
210 - 80 = 130 km the Van is from A
:
Find how far the Van has traveled from B at 3:30
300 - 130 = 170 km
:
Find the speed of the Van; speed = dist/time
170/3.5 = 48.6 km/hr
;
Find how long the Van takes to travel 130 km to A, at 48.6 km/hr
130/48.6 = 2.73 hrs = 2 + .73(60) = 2:44 hrs & minutes
:
Find the time it arrives at A
3:30 + 2:44 = 6:14 PM